# Mamba-in-Mamba: Centralized Mamba-Cross-Scan in Tokenized Mamba Model for Hyperspectral Image Classification

### v3

@@ III-B2 MCS Type-2
$\mathbf{S}^{(i,j)}_{2,\,f}$ $=\Big{[}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbf{x}^{(i-2,j-2)}},\mathbf{x}^{(i-2,j-1)},\mathbf{x}^{(i-2,j)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (17a)
$\mathbf{S}^{(i,j)}_{2,\,b}$ $=\Big{[}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{x}^{(i+2,j+2)}},\mathbf{x}^{(i+2,j+1)},\mathbf{x}^{(i+2,j)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top}.$ (17b)
@@
III-B3 MCS Type-3
Different from the type-1 and type-2 which starts from $\mathbf{x}^{(i-2,\,j-2)}$ and $\mathbf{x}^{(i+2,\,j+2)}$ positions, the type-3 cross Mamba scan is starting from $\mathbf{x}^{(i-2,\,j+2)}$ and $\mathbf{x}^{(i-2,\,j+2)}$ with a horizontal manner, shown in Fig. 5(c). The formulated two sub-sequences are listed below:
$\mathbf{S}^{(i,j)}_{3,\,f}$ $=\Big{[}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbf{x}^{(i-2,j+2)}},\mathbf{x}^{(i-1,j+2)},\mathbf{x}^{(i,j+2)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (18a)
$\mathbf{S}^{(i,j)}_{3,\,b}$ $=\Big{[}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{x}^{(i+2,j-2)}},\mathbf{x}^{(i+1,j-2)},\mathbf{x}^{(i,j-2)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top}.$ (18b)
III-B4 MCS Type-4
@@
$\mathbf{S}^{(i,j)}_{4,\,f}$ $=\Big{[}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbf{x}^{(i-2,j+2)}},\mathbf{x}^{(i-2,j+1)},\mathbf{x}^{(i-2,j)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (19a)
$\mathbf{S}^{(i,j)}_{4,\,b}$ $=\Big{[}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{x}^{(i+2,j+2)}},\mathbf{x}^{(i+2,j+1)},\mathbf{x}^{(i+2,j)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (19b)
@@
III-C Tokenized Mamba (T-Mamba)
@@
III-C1 Whole Flowchart of T-Mamba
@@
Subsequently, these processed sub-sequences are input to the Mamba model as described in Section. II:
@@
III-C2 Gaussian Decay Mask (GDM)
@@
III-C3 Cross Scan Merge
@@
III-C4 Semantic Token Learner (STL)
@@
III-C5 Semantic Token Fuser (STF)
@@
Figure 9: Four datasets, (a) Indian Pines (IP), (b) Pavia University (PU), (c) Houston 2013 (HU2013), and (d) WHU-Hi-HongHu (HongHu) are used for experiments. The false-color images, fixed and disjoint training sets and testing sets are shown above.
TABLE I: Description of Land-Cover Types and Number of Train-Test Samples for Four Datasets
@@ 22                 Tree 25 4,015
TABLE II: The Hyperparameter settings and size of notations for Four datasets.
@@ III-D Weighted MCS Fusion (WMF)
TABLE III: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE INDIAN PINES DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
No.     Color.     vanilla ViT[18]     SpeFormer[19]     MAEST[21]     HiT[22]     SSFTT[26]     3DViT[27]     HUSST[28]     ViM[52]     MiM (Ours)
@@
TABLE IV: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE PAVIA UNIVERSITY DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
No.     Color.     vanilla ViT[18]     SpeFormer[19]     MAEST[21]     HiT[22]     SSFTT[26]     3DViT[27]     HUSST[28]     ViM[52]     MiM (Ours)
@@
III-E Multi-Scale Loss Design
@@
IV Experimental results and analysis
In this section, we first describe four well-known HSI datasets. We then outline the implementation details and the methods used for comparison. Finally, we conduct extensive experiments to evaluate the performance of the proposed method both quantitatively and qualitatively.
IV-A Datasets Description
Four HSI datasets are illustrated in Fig. 9, along with their corresponding false-color images, training samples and testing samples. The numbers of training and testing samples for each class within these datasets are listed in Table I.
@@ IV-A Datasets Description
IV-B Experimental Evaluation Indicators
To quantitatively assess the effectiveness of our proposed method and compare it with other methods, we rely on four key performance indicators: overall accuracy (OA), average accuracy (AA), the kappa coefficient (kappa), and accuracy within each class. These indicators provide a comprehensive view of the classification performance, with higher values indicating superior classification results.
@@ IV-B Experimental Evaluation Indicators
TABLE V: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE HOUSTON 2013 DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
No.     Color.     vanilla ViT[18]     SpeFormer[19]     MAEST[21]     HiT[22]     SSFTT[26]     3DViT[27]     HUSST[28]     ViM[52]     MiM (Ours)
@@
IV-C Brief Description and Settings of Compared Methods
To evaluate the effectiveness of our proposed method, we conducted comparative analyses against a range of baseline and state-of-the-art models. These included vanilla ViT (ViT) [18], SpeFormer [19], MAEST [21], HiT [22], SSFTT [26], 3DViT [27], HUSST [28], and ViM [52]. All models were employed as patch-wise classifiers, with parameter settings adhering to their respective references, especially the initial patch size for different datasets. The comparison is grounded in the same number of the training samples without data augmentation.
@@ IV-C Brief Description and Settings of Compared Methods
TABLE VI: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE WHU-Hi-HongHu DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
No.     Color.     vanilla ViT[18]     SpeFormer[19]     MAEST[21]     HiT[22]     SSFTT[26]     3DViT[27]     HUSST[28]     ViM[52]     MiM (Ours)
@@
IV-D Experimental Results
IV-D1 Quantitative Results and Classification Maps
Tables III, IV, V, and VI present average classification metrics under 5 times of running, such as OA, AA, Kappa coefficient, and class-wise accuracy for the IP, PU, HU2013, and HongHu datasets, respectively. Corresponding classification maps are displayed in Figs. 10, 11, 12, and 13.
@@ IV-D1 Quantitative Results and Classification Maps
TABLE VII: TOTAL RUNNING TIME (MINUTES) AND THE NUMBER OF PARAMETERS REQUIRED BY DIFFERENT METHODS ON THE INDIAN PINES DATASET.
vanilla ViT[18]     SpeFormer[19]     MAEST[21]     HiT[22]     SSFTT[26]     3DViT[27]     HUSST[28]     ViM[52]     MiM (Ours)
@@
IV-D2 Learned Feature Visualizations by T-SNE
@@
IV-D3 Comparison of Computational Costs
To evaluate computational efficiency across all methods, we conducted a detailed statistical analysis, summarized in Table VII. This analysis adheres to previously described experimental setups and includes metrics such as total training time (minutes), testing time (seconds), and model size (number of parameters) for the IP dataset.
@@ IV-D3 Comparison of Computational Costs
IV-E Other Analysis and Ablation Study
@@
TABLE VIII: IMPACT OF DIFFERENT INSERTABLE COMPONENTS IN T-MAMBA ENCODER ON THE CLASSIFICATION RESULTS (%). THE BEST ONE IS HIGHLIGHTED AS RED.
@@
IV-E1 Ablation on Proposed Different Components
In the T-Mamba encoder, we integrated several key components: the STL, GDM, and STF. To assess their effectiveness, we conducted an ablation study, the results of which are summarized in Table VIII. The ablation study details how each component was tested. In scenarios where STL and STF were not used, we substituted them with general average pooling and simple summation, respectively, to gauge their relative performance.
@@ IV-E1 Ablation on Proposed Different Components
TABLE IX: IMPACT OF DIFFERENT INITIAL PATCH SIZE FOR THE OVERALL ACCURACY (%) ON FOUR DATASETS. THE BEST ONE IS HIGHLIGHTED AS RED.
@@
IV-E2 Analysis on Initial Patch Size
@@
As shown in Table IX, we evaluated a range of initial patch sizes (3, 5, 7, 9, 11, 13, 15). The results, specifically OA for four datasets, are detailed in the same table.
@@
IV-E3 Investigation of Scan Design
@@
TABLE X: THE COMPARISON OF USING DIFFERENT SCAN DESIGNS FOR OA (%). THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
@@
We tested these scanning methods to evaluate their performance, with results summarized in Table X. Each scan design was tested using one to four types. The most effective results were achieved using the Mamba-scan with four directional types, confirming that a continuous sequence arrangement enhances content recognition. This finding is further supported by comparing the Diagonal-scan and the Zig-zag-scan, where the latter showed slight improvements due to its semi-continuous positioning. The commonly used Raster-scan, despite its prevalence in image transformation tasks, tends to disrupt spatial continuity, particularly when spatial context is significant in RS land-cover analysis, resulting in inferior performance compared to the Mamba-scan.
@@
IV-E4 Exploring the Feature Fusion Strategy in T-Mamba
We developed our bi-directional feature fusion strategy in T-Mamba encoder with two ways, as shown in Fig. 16: 1) Post-Merge strategy, and 2) Pre-Merge strategy. To evaluate the effectiveness of these two strategies, we conducted experiments across four datasets. The results of these experiments are detailed in Table XI.
TABLE XI: IMPACT OF DIFFERENT FEATURE FUSION STRATEGY IN T-MAMBA ENCODER FOR THE OA (%) ON FOUR DATASETS. THE BEST ONE IS HIGHLIGHTED AS RED.

### v4

@@ 3.2.2 MCS Type-2
$\mathbf{S}^{(i,j)}_{2,\,f}$ $=\Big{[}{\color[rgb]{0,1,0}\mathbf{x}^{(i-2,j-2)}},\mathbf{x}^{(i-2,j-1)},\mathbf{x}^{(i-2,j)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (17a)
$\mathbf{S}^{(i,j)}_{2,\,b}$ $=\Big{[}{\color[rgb]{0,0,1}\mathbf{x}^{(i+2,j+2)}},\mathbf{x}^{(i+2,j+1)},\mathbf{x}^{(i+2,j)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top}.$ (17b)
@@
3.2.3 MCS Type-3
Different from the type-1 and type-2 which starts from two positions of $\mathbf{x}^{(i-2,\,j-2)}$ and $\mathbf{x}^{(i+2,\,j+2)}$, the type-3 cross Mamba scan is starting from $\mathbf{x}^{(i-2,\,j+2)}$ and $\mathbf{x}^{(i-2,\,j+2)}$ with a horizontal manner, shown in Fig. 5(c). The formulated two sub-sequences are listed below:
$\mathbf{S}^{(i,j)}_{3,\,f}$ $=\Big{[}{\color[rgb]{0,1,0}\mathbf{x}^{(i-2,j+2)}},\mathbf{x}^{(i-1,j+2)},\mathbf{x}^{(i,j+2)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (18a)
$\mathbf{S}^{(i,j)}_{3,\,b}$ $=\Big{[}{\color[rgb]{0,0,1}\mathbf{x}^{(i+2,j-2)}},\mathbf{x}^{(i+1,j-2)},\mathbf{x}^{(i,j-2)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top}.$ (18b)
3.2.4 MCS Type-4
@@
$\mathbf{S}^{(i,j)}_{4,\,f}$ $=\Big{[}{\color[rgb]{0,1,0}\mathbf{x}^{(i-2,j+2)}},\mathbf{x}^{(i-2,j+1)},\mathbf{x}^{(i-2,j)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (19a)
$\mathbf{S}^{(i,j)}_{4,\,b}$ $=\Big{[}{\color[rgb]{0,0,1}\mathbf{x}^{(i+2,j+2)}},\mathbf{x}^{(i+2,j+1)},\mathbf{x}^{(i+2,j)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}\Big{]}^{\top},$ (19b)
@@
3.3 Tokenized Mamba (T-Mamba)
@@
3.3.1 Whole Flowchart of T-Mamba
@@
Subsequently, these processed sub-sequences are input to the Mamba (S6) model as described in Section. 2:
@@
3.3.2 Gaussian Decay Mask (GDM)
@@
3.3.3 Cross Scan Merge
@@
3.3.4 Semantic Token Learner (STL)
@@
3.3.5 Semantic Token Fuser (STF)
@@
Algorithm 1 Summary of procedure for the proposed method
2:  Output: Predicted labels of the testing samples.
4:  Procedure:
9:  MiM model:
10:  Apply four types of MCS on $\textbf{X}^{(i,\,j)}$, getting sequences, $\mathbf{S}_{1}^{(i,\,j)}$, $\mathbf{S}_{2}^{(i,\,j)}$, $\mathbf{S}_{3}^{(i,\,j)}$, and $\mathbf{S}_{4}^{(i,\,j)}$ (Eqs. 14a-19a).
11:  for $m=1$ to $4$ in $\mathbf{S}_{m}^{(i,\,j)}$ do
13:     T-Mamba encoder:
14:     for Both forward pass $\mathbf{s}_{f}$ and backward pass $\mathbf{s}_{b}$ do
21:        Activate $\mathbf{o}_{m}^{(i,\,j)}$ using the Tanh function.
22:     end for
23:  end for
Table 1: Description of land-cover types and number of train-test samples for four datasets
@@ 22 Tree 25 4,015
Table 2: The hyperparameter settings and size of notations for four datasets.
@@ 3.4 Weighted MCS Fusion (WMF)
Table 3: Different classification methods in terms of OA, AA, and Kappa as well as the accuracies for each class on the Indian Pines dataset. The best one is highlighted as red. The second best is highlighted as blue.
No. Color. vanilla ViT SpeFormer MAEST HiT SSFTT 3DViT HUSST ViM MiM (Ours)
@@
Table 4: Different classification methods in terms of OA, AA, and Kappa as well as the accuracies for each class on the Pavia University dataset. The best one is highlighted as red. The second best is highlighted as blue.
No. Color. vanilla ViT SpeFormer MAEST HiT SSFTT 3DViT HUSST ViM MiM (Ours)
@@
3.5 Multi-Scale Loss Design
@@
Figure 9: Four experimental datasets, (a) Indian Pines (IP), (b) Pavia University (PU), (c) Houston 2013 (HU2013), and (d) WHU-Hi-HongHu (HongHu) are used for experiments. The false-color images, fixed and disjoint training sets and testing sets are shown above.
4 Experimental results and analysis
In this section, we first describe four well-known HSI datasets. We then outline the implementation details and the methods used for comparison. Finally, we conduct extensive experiments to evaluate the performance of the proposed method both quantitatively and qualitatively. Due to the paper formatting, all evaluation tables are placed at end of the article.
4.1 Datasets Description
Four HSI datasets are illustrated in Fig. 9, along with their corresponding false-color images, training and testing samples. The numbers of training and testing samples for each class within these datasets are listed in Table 1.
@@ 4.1 Datasets Description
4.2 Experimental Evaluation Indicators
To quantitatively assess the effectiveness of our proposed method and compare it with other methods, we rely on four key performance indicators: overall accuracy (OA), average accuracy (AA), the kappa coefficient (kappa), and accuracy within each class. They provide a comprehensive view of the classification performance, with higher values indicating superior classification results.
@@ 4.2 Experimental Evaluation Indicators
Table 5: Different classification methods in terms of OA, AA, and Kappa as well as the accuracies for each class on the Houston 2013 dataset. The best one is highlighted as red. The second best is highlighted as blue.
No. Color. vanilla ViT SpeFormer MAEST HiT SSFTT 3DViT HUSST ViM MiM (Ours)
@@
4.3 Brief Description and Settings of Compared Methods
We conducted comparative analyses against a range of baseline and state-of-the-art models. These included vanilla ViT (ViT) [18], SpeFormer [19], MAEST [21], HiT [22], SSFTT [26], 3DViT [27], HUSST [28], and ViM [51]. All models were employed as patch-wise classifiers, with parameter settings adhering to their respective references, especially the initial patch size for different datasets. The comparison is grounded in the same number of the training samples without data augmentation.
@@ 4.3 Brief Description and Settings of Compared Methods
Table 6: Different classification methods in terms of OA, AA, and Kappa as well as the accuracies for each class on the WHU-Hi-HongHu dataset. The best one is highlighted as red. The second best is highlighted as blue.
No. Color. vanilla ViT SpeFormer MAEST HiT SSFTT 3DViT HUSST ViM MiM (Ours)
@@
4.4 Experimental Results
4.4.1 Quantitative Results and Classification Maps
Tables 3, 4, 5, and 6 present average classification metrics under 5 times of running, such as OA, AA, Kappa coefficient, and class-wise accuracy for the IP, PU, HU2013, and HongHu datasets, respectively. Corresponding classification maps are displayed in Figs. 10, 11, 12, and 13.
@@ 4.4.1 Quantitative Results and Classification Maps
Table 7: Total running time (minutes) and the number of parameters required by different methods on the Indian Pines dataset.
vanilla ViT SpeFormer MAEST HiT SSFTT 3DViT HUSST ViM MiM (Ours)
@@
Key differences in classification performance are highlighted and marked in Figs. 10, and 11 for detailed comparisons.
@@
4.4.2 Learned Feature Visualizations by T-SNE
@@
4.4.3 Comparison of Computational Costs
To evaluate computational efficiency across all methods, we conducted a detailed statistical analysis, summarized in Table 7. This analysis adheres to previously described experimental setups and includes metrics such as total training time (minutes), testing time (seconds), and model size (number of parameters) for the IP dataset.
@@ 4.4.3 Comparison of Computational Costs
4.5 Other Analysis and Ablation Study
@@
Table 8: Impact of different insertable components in T-Mamba encoder on the overall classification results (%). The best one is highlighted as red.
@@
4.5.1 Ablation on Proposed Different Components
In the T-Mamba encoder, we integrated several key components: the STL, GDM, and STF. To assess their effectiveness, we conducted an ablation study, the results of which are summarized in Table 8. The ablation study details how each component was tested. In scenarios where STL and STF were not used, we substituted them with general average pooling and simple summation, respectively, to gauge their relative performance.
@@ 4.5.1 Ablation on Proposed Different Components
Table 9: Impact of different initial patch size for the overall accuracy (%) on four datasets. The best one is highlighted as red.
@@
4.5.2 Analysis on Initial Patch Size
@@
As shown in Table 9, we evaluated a range of initial patch sizes (3, 5, 7, 9, 11, 13, 15). The results, specifically OA for four datasets, are detailed in the same table.
@@
4.5.3 Investigation and Comparison on Scan Design
@@
Table 10: The comparison of using different scan designs for OA (%). The best one is highlighted as red. The second best is highlighted as blue.
@@
We tested these scanning methods to evaluate their performance, with results summarized in Table 10. Each scan design was tested using one to four types. The most effective results were achieved using the Mamba-scan with four directional types, confirming that a continuous sequence arrangement enhances content recognition. This finding is further supported by comparing the Diagonal-scan and the Zig-zag-scan, where the latter showed slight improvements due to its semi-continuous positioning. The commonly used Raster-scan, despite its prevalence in image transformation tasks, tends to disrupt spatial continuity, particularly when spatial context is significant in RS land-cover analysis, resulting in inferior performance compared to the Mamba-scan.
@@
4.5.4 Exploring the Feature Fusion Strategy in T-Mamba
We developed our bi-directional feature fusion strategy in T-Mamba encoder with two ways, as shown in Fig. 16: 1) Post-Merge strategy, and 2) Pre-Merge strategy. To evaluate the effectiveness of these two strategies, we conducted experiments across four datasets. The results of these experiments are detailed in Table 11.
Table 11: Impact of different feature fusion strategy in T-Mamba encoder for the OA (%) on four datasets. The best one is highlighted as red.